OpenStudy (anonymous):

Find S10 for a geometric series with first term 10 and common ratio 4.

OpenStudy (curry):

3495250 but im not sure

OpenStudy (anonymous):

s10= 10(4^10-1)/(4-1)=10(1048575)/3=3495250

OpenStudy (curry):

yep

OpenStudy (anonymous):

Sn = a (r^n - 1) / (r-1) 3495250

OpenStudy (anonymous):

you are correct currey:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks guys :)